# Fully quantum algorithm for mesoscale fluid simulations with application to partial differential equations

### v3

@@ IV.2 Measurement-avoiding QLBM
An issue with QLGA is that the quantum state collapses due to measurements performed during each time step, requiring re-encoding. However, we should allow the qubits to explore more complex states to achieve truly quantum computation. Furthermore, measurement on superconducting quantum computers can typically take an order of magnitude longer than gate times, adding significant time to the algorithm [37]. Thus, the goal is to eliminate repeated initialization of the quantum state and repeated measurements, which destroy the quantum superposition and entanglement.
@@
To this end, our algorithm adds a phase-correction step, approximating the qubit relative phases and subtracting them at the end of each time step, delaying the need for measurement until the end of the computation. This can be achieved using the quantum phase estimation (QPE) algorithm as a sub-routine to estimate the eigen-phases of the encoding and collision unitaries [38]. These eigen-phases are then used to classically estimate the qubit relative phases, subtracted at the end of each time step. The accuracy of the QPE depends on the number of ancillary qubits it uses. Thus, we use the iterative phase estimation (IPE) algorithm, which only requires a single auxiliary qubit and evaluates the phase bit by bit through a repetitive process [39]. The accuracy of the IPE algorithm is determined by the number of iterations instead of the number of ancillary qubits. Therefore, they are of great importance for near-term quantum computing. The IPE algorithm still requires measurement to estimate the phase. However, only one qubit, i.e., the auxiliary qubit, is measured, and not all qubits are in the quantum register.
Even though the IPE algorithm requires measurements, the measurements required by IPE are minimal compared to those required to read the quantum circuit state and the subsequent complexity of re-encoding it. Due to this, the measurements required by IPE enable the overall technique to avoid excessive measurement operations.
(a) First two iterations of IPE algorithm
(b) Measurement-avoiding QLBM circuit
@@
Figure 5: (a) IPE circuit used to estimate the qubits relative phase $\alpha_{i}$ and (b) Measurement-avoiding QLBM circuit solving the diffusion equation.
@@
In the context of QLBM, we apply the IPE algorithm to every lattice site, i.e., to a two-qubit unitary for the D1Q2 scheme. We set the unitary $U$ to be the product of the encoding and the collision operations since the streaming step only swaps the qubits and does not change their phases. For a total number of iterations $m$, the IPE algorithm estimates the phase $\varphi=0.\varphi_{1}...\varphi_{m}$, where $\varphi_{k}$ are the phase bits ($1\leq k\leq m$). At each iteration, $k$, the IPE algorithm applies the unitary $U$ $2^{m-k}$ times controlled on the auxiliary qubit. Due to phase kickback, the relative phase of the auxiliary qubit becomes $\exp(i2\pi 0.\varphi_{m-k+1}\varphi_{m-k+2}...\varphi_{m})$. Thus, to estimate the phase bit $\varphi_{m-k+1}$, a phase correction of $-2\pi(\varphi_{m-k+2}/2^{2}+...+\varphi_{m}/2^{m})$ is performed, controlled on the measurement outcome of the previously estimated phase bits $\varphi_{m-k+2},...,\varphi_{m}$. The phase bit $\varphi_{m-k+1}$ is then estimated by measuring the auxiliary qubit in the $x$-basis. The first two iterations of the IPE, which estimates the two least significant bits of the phase $\varphi_{m}$ and $\varphi_{m-1}$, is illustrated in fig. 5 (a). One time step of the measurement-avoiding QLBM circuit is shown in the circuit of fig. 5 (b), which consists of collision, streaming, and phase correction operations.
@@
Figure 6 presents the numerical simulation of our variant using the IPE algorithm to solve the 1D diffusion equation. The simulation uses the D1Q2 scheme with 26 qubits and runs on the statevector simulator. The spatial domain consists of $N_{s}=13$ lattice sites, and results are shown after $6$ time steps. A good agreement is found between the quantum solution and the classical one. This variant improves the efficiency of the QLGA algorithm as the encoding is used only once at the beginning of the computation to prepare the quantum states, and measurement is only performed at the end of the computation.
General quantum devices have expensive measurement steps, especially in the context of time-stepping problems. Avoiding such measurement is thus a prerequisite of a quantum algorithm for any quantum device. This also leads to fewer state-preparations steps, which are computationally expensive in comparison the the other algorithmic considerations.
However, this algorithm is hybrid since the qubits phases are classically computed from the eigenvalues estimated using the IPE. Regarding complexity, having only one ancillary qubit comes with the cost of $O(2^{m-1})$ gates, where $m$ is the number of IPE iterations related to the precision of the estimated phases. Given unbiased assumptions, the finite sampling error of iterative phase estimation scales proportional to $N^{-1/2}$. Tomographic estimation requires estimating a state with multiple parameters while undoing phase estimation requires only one estimation. A qubit state is multi-parametric, so subtracting the phase with a single estimation is more efficient.
@@
Figure 6: Measurement-avoiding QLBM (present work) solution to the diffusion equation using the IPE algorithm and its comparison to classical results for the same method and resolution.
@@ V Results
(a) RMSE compared to the classical solution using $10^{4}$ shots
(b) RMSE compared to the classical solution using $18$ qubits
Figure 7: RMSE of the quantum solution solving the diffusion equation compared to the classical one. Qiskit (solid lines, ——) statevector and MPS simulators and XACC simulators (dashed lines, – – –) qpp and exatn-mps are shown. The colors correspond to (blue) statevector-like and (orange) matrix product state simulators.
@@
where $N_{s}$ is the number of lattice sites of the quantum solution compared to the classical one as a function of the number of shots (samples from the underlying density functions) and qubits. The errors are reported from simulations using $10^{4}$ shots in fig. 7 (a) and using $18$ qubits in fig. 7 (b).
These figures demonstrate the scalability of this method. The RSME for the algorithm follows the trend of decreasing quadratically with the number of shots, as shown in fig. 7 (b). An RMSE of $0.01$ occurs at less than $10^{4}$ shots. This trend occurs when the algorithm is run on quantum hardware, as the state is read by taking samples from the resulting distribution. The number of qubits does not affect the RMSE greatly, as shown in fig. 7 a, and the technique can scale well to more lattice sites with a linear increase in qubits.
The MPS simulator uses a local representation in a factorized form of tensor products [40]. The MPS method ensures that the overall structure remains small as long as the circuit has a low degree of entanglement, enabling more simulated qubits. To demonstrate this, we run the QLBM circuit of fig. 3 (a) on the Qiskit MPS simulator for a resolution of $51$ lattice sites corresponding to $102$ qubits using $10^{6}$ shots. Results after $10$ time steps are shown in fig. 8 and demonstrate good agreement with the classical solution.
@@ VI Conclusion and Discussion
In this work, we present a revised quantum lattice gas algorithm that can be used to simulate fluid flows. The algorithm of Yepez [41] was designed for Type-II quantum computers that are no longer used. The presented algorithm uses a full-quantum implementation of the streaming step to eliminate classical communication between lattice sites. In addition, it avoids repetitive measurements and encoding operations by estimating the qubit relative phases and subtracting them at the end of each time step. This eliminates the need for state re-preparation between time steps and delays the need for measurement to the end of the computation. State re-preparation and measurements can scale exponentially in modern computers, and our work on measurement avoidance is essential for an efficient adaptation of the QLGA to modern hardware. The algorithm is tested to solve canonical PDEs: the diffusion and Burgers’ equations. We explored different quantum simulators available on Qiskit and XACC quantum computing frameworks and compared their performance and wall-clock time. Matrix product state simulators are most efficient for the presented quantum circuits, as they have low and local entanglement. These strategic algorithm improvements enable the execution of the quantum lattice gas algorithm on quantum hardware, a subject of future work.
An important limitation of the current algorithm is the size of the required quantum register, which scales linearly with the number of lattice sites. For example, the number of required qubits in a two-dimensional $M\times M$ lattice is $5M^{2}$ using the D2Q5 scheme. This requirement constrains the ability to extend this algorithm to higher dimensions due to the limited number of qubits available on current quantum hardware. Thus, one needs to use more appropriate encoding schemes, such as one that scales logarithmically with the number of lattice sites, like amplitude-based encoding. For example, Budinski [22] implements the streaming step via a quantum random walk, similar to the collisionless Boltzmann equation solved on a quantum computer of Todorova and Steijl [21]. Indeed, in principle, one could design an algorithm that uses a quantum random walk for both the collision and streaming operations, which we will investigate in future work.
@@ Acknowledgements
We thank Dr. Jeffrey Yepez for numerous fruitful discussions of this work. S.H.B. acknowledges support from the Georgia Tech Quantum Alliance and a Georgia Tech Seed Grant. E.F.D. acknowledges DOE ASCR funding under the Quantum Computing Application Teams program, FWP number ERKJ347. This work used Bridges2 at the Pittsburgh Supercomputing Center through allocation PHY210084 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.
@@ Appendix A D1Q3 Implementation
The implementations of the QLBM mentioned in section II and section III used the D1Q2 scheme, which uses two moving particles per lattice site. This scheme shows checker-board pathologies when using a sharp function as the initial condition because it simulates two independent sub-lattice [26]. Yepez [36] presents a cure for this, allowing only one qubit to move while the other remains stationary. However, this requires two collision and streaming procedures per time step. We show that a quantum D1Q3 scheme that adds a stationary qubit to each lattice site solves this problem. A D1Q3 scheme has three particle distributions across 3 links: left, right, and stationary. This means some fictitious particles remain stationary in a timestep, smoothing the curve. Figure 9 compares the quantum solution of the diffusion equation using the D1Q2 scheme with the D1Q3 scheme for a delta function as the initial condition. These simulations were performed using the Qiskit matrix product state simulator, $45$ qubits for the D1Q3 scheme, and the sampling procedure used $10^{6}$ shots. The spatial domain consists of $N_{s}=15$ lattice sites, and results are shown after $10$ time steps. The result shows the agreement of the quantum D1Q3 solution with the classical one and how the D1Q3 version remedies the checker-board effect. We note that the D1Q3 scheme does not affect algorithmic complexity and can employ measurement avoidance analogous to the D1Q2 scheme presented in the main manuscript body.
Figure 9: QLBM solutions to the diffusion equation using the D1Q3 scheme. “Quantum - classical streaming” is associated with the classical streaming algorithm discussed in previous work [26], “Fully quantum” refers to the present algorithm, and “Classical” is the classical implementation of the lattice methods as labeled.
@@ Appendix B Simulator Results
In fig. 10 we show the results for different quantum simulator implementations of our algorithm and its circuit. Simulators include statevector-based and tensor-network types that use the matrix product state (MPS) method [42]. Two different simulation frameworks are tested: Qiskit and XACC [43]. Statevector and MPS simulators are part of the Qiskit Aer backend. The XACC qpp (Quantum++) simulator is based on a C++ general-purpose quantum computing library [44]. XACC exatn-mps simulator is a part of the TNQVM (tensor-network quantum virtual machine) simulation backend and uses a noiseless, matrix product state wave function decomposition for the quantum circuit [45]. The wall-clock times of fig. 7 (c) were obtained via simulation on an AMD EPYC 7742 (64-core) processor.

### v4

@@ IV.2 Restless QLBM
An issue with QLGA is that the quantum state collapses due to measurements performed during each time step, requiring re-encoding. However, we should allow the qubits to explore more complex states to achieve truly quantum computation. Furthermore, measurement on superconducting quantum computers can typically take an order of magnitude longer than gate times, adding significant time to the algorithm [39]. Thus, the goal is to eliminate repeated initialization of the quantum state and repeated measurements.
@@
To this end, our algorithm adds a phase-correction step, approximating the qubit relative phases and subtracting them at the end of each time step, delaying the need for measurement until the end of the computation. Phase correction is achieved via quantum phase estimation (QPE) as a sub-routine to estimate the eigen-phases of the encoding and collision unitaries [30]. These eigen-phases are then used to classically estimate the qubit relative phases, subtracted at the end of each time step. The accuracy of the QPE depends on the number of ancillary qubits it uses. Thus, we use the iterative phase estimation (IPE) algorithm, which only requires a single auxiliary qubit and evaluates the phase bit by bit through a repetitive process [40]. The accuracy of the IPE algorithm is determined by the number of iterations instead of the number of ancillary qubits. Therefore, they are of great importance for near-term quantum computing. The IPE algorithm still requires measurement to estimate the phase. However, only one qubit, i.e., the auxiliary qubit, is measured, and not all qubits are in the quantum register.
Although the phase estimation algorithm could help avoid measurements at the end of each time step and re-preparation in the next time step, it introduces extra $O(n)$ mid-circuit measurements in each time step. This overhead balances off the save for measurement. What’s left is the resource reduction for state re-preparation with $O(T)$ circuit depth, where $T$ is the number of time steps.
Refer to caption Refer to caption
(a) (b)
@@
Figure 5: (a) IPE circuit used to estimate the qubits relative phase $\alpha_{i}$ (first two iterations) and (b) Restless QLBM circuit solving the diffusion equation.
@@
In the context of QLBM, we apply the IPE algorithm to every lattice site, i.e., to a two-qubit unitary for the D1Q2 scheme. We set the unitary $U$ to be the product of the encoding and the collision operations since the streaming step only swaps the qubits and does not change their phases. For a total number of iterations $m$, the IPE algorithm estimates the best $m$ bit approximation to $\phi$ with $\varphi=0.\varphi_{1}...\varphi_{m}=\sum_{k=1}^{m}\varphi_{k}/2^{k}$, where $\varphi_{k}$ are the phase bits ($1\leq k\leq m$) in binary expansion. At each iteration, $k$, the IPE algorithm applies the unitary $U$ $2^{m-k}$ times controlled on the auxiliary qubit. Due to phase kickback, the relative phase of the auxiliary qubit becomes $\exp(i2\pi 0.\varphi_{m-k+1}\varphi_{m-k+2}...\varphi_{m})$. Thus, to estimate the phase bit $\varphi_{m-k+1}$, a phase correction of $-2\pi(\varphi_{m-k+2}/2^{2}+...+\varphi_{m}/2^{m})$ is performed, controlled on the measurement outcome of the previously estimated phase bits $\varphi_{m-k+2},\dots,\varphi_{m}$. The phase bit $\varphi_{m-k+1}$ is then estimated by measuring the auxiliary qubit in the $x$-basis. The first two iterations of the IPE algorithm, which estimates the two least significant bits of the phase $\varphi_{m}$ and $\varphi_{m-1}$, is illustrated in fig. 5 (a). One time step of the restless QLBM circuit is shown in the circuit of fig. 5 (b), which consists of collision, streaming, and phase correction operations.
@@
Figure 6 presents the numerical simulation of our variant using the IPE algorithm to solve the 1D diffusion equation. The simulation uses the D1Q2 scheme with 26 qubits and runs on the statevector simulator. The spatial domain consists of $N_{s}=13$ lattice sites, and results are shown after $6$ time steps. The quantum solution and the classical one agree well, with the root-mean-square error between them below the finite sampling threshold. This variant improves the efficiency of the QLGA algorithm as the encoding is used only once at the beginning of the computation to prepare the quantum states, and measurement is only performed at the end of the computation.
Refer to caption Refer to caption
(a) (b)
@@
Figure 6: (a) Restless QLBM (present work) solution to the diffusion equation using the IPE algorithm with $m=14$ iterations, $N=10^{5}$ shots, and its comparison to classical results for the same method and resolution. (b) Comparing different number of iterations $m$ impact the overall algorithmic error.
The number of iterations can be determined by a desired accuracy $1/2^{q}$ with probability of success at least $p$
$m=q+\bigg{\lceil}\log_{2}\bigg{(}2+\frac{1}{2(1-p)}\bigg{)}\bigg{\rceil}.$ (12)
[…]
Regarding complexity, having only one ancillary qubit comes with the cost of $O(2^{m-1})$ gates, where $m$ is the number of IPE iterations related to the precision of the estimated phases. Given unbiased assumptions, the finite sampling error of iterative phase estimation scales proportional to $1/\sqrt{N}$, where $N$ is the shot number in quantum measurement. A qubit state is multi-parametric, so subtracting the phase with a single estimation is more efficient.
@@ V Results
where $N_{s}$ is the number of lattice sites of the quantum solution compared to the classical one as a function of the number of shots (samples from the underlying density functions) and qubits. The errors are reported from simulations as a function of the number of lattice sites in fig. 7 (a) and as a function of the number of shots in fig. 7 (b).
Refer to caption Refer to caption
(a) (b)
Figure 7: RMSE of the quantum solution solving the diffusion equation compared to the classical one. (a) $10^{4}$ shots and (b) 18 qubits. Qiskit (solid lines, ——) statevector and MPS simulators and XACC simulators (dashed lines, – – –) qpp and exatn-mps are shown. The colors correspond to (blue) statevector-like and (orange) matrix product state simulators. Notably, each lattice site contains $2$ qubits for a D1Q2 scheme.
These figures demonstrate the scalability of this method. The RMSE for the algorithm follows the trend of decreasing quadratically with the number of shots, as shown in fig. 7 (b). An RMSE of $0.01$ occurs at less than $10^{4}$ shots. This accuracy level is proportional to $1/\sqrt{N}$ due to finite sampling of the quantum state, where $N$ is the shot number. The number of qubits does not affect the RMSE greatly, as shown in fig. 7 (a), and the technique can scale well to more lattice sites with a linear increase in qubits.
The MPS simulator uses a local representation in a factorized form of tensor products [41]. MPS ensures that the overall structure remains small as long as the circuit has a low degree of entanglement, enabling more simulated qubits. To demonstrate this, we run the QLBM circuit of fig. 3 (a) on the Qiskit MPS simulator for a resolution of $51$ lattice sites corresponding to $102$ qubits using $10^{6}$ shots. Results after $10$ time steps are shown in fig. 8 and demonstrate good agreement with the classical solution.
@@ VI Conclusion and Discussion
In this work, we present a revised quantum lattice gas algorithm that can be used to simulate fluid flows. The algorithm of Yepez [42] was designed for Type-II quantum computers that are no longer used. The presented algorithm uses a full-quantum implementation of the streaming step to reduce classical communication between lattice sites. In addition, quantum streaming optimizes repetitive encoding operations by estimating the qubit relative phases and subtracting them at the end of each time step. State re-preparation and measurements can scale exponentially in modern computers, and our work on state preparation optimization is essential for efficiently adapting the QLGA to modern hardware. The algorithm is tested to solve canonical PDEs: the diffusion and Burgers’ equations. We explored different quantum simulators available on Qiskit and XACC quantum computing frameworks and compared their performance and wall-clock time. Matrix product state simulators are most efficient for the presented quantum circuits, as they have low and local entanglement. These strategic algorithm improvements enable the execution of the quantum lattice gas algorithm on quantum hardware, a subject of future work.
An important limitation of the current algorithm is the size of the required quantum register, which scales linearly with the number of lattice sites. For example, the number of required qubits in a two-dimensional $M\times M$ lattice is $5M^{2}$ using the D2Q5 scheme. This requirement constrains the ability to extend this algorithm to higher dimensions due to the limited number of qubits available on current quantum hardware. Thus, more appropriate encoding schemes, such as amplitude-based encoding, are needed to scale logarithmically with the number of lattice sites. For example, Budinski [22] implements the streaming step via a quantum random walk, similar to the collisionless Boltzmann equation solved on a quantum computer of Todorova and Steijl [21]. Indeed, in principle, one could design an algorithm that uses a quantum random walk for both the collision and streaming operations, which we will investigate in future work.
@@ Acknowledgements
We thank Dr. Jeffrey Yepez for numerous fruitful discussions of this work. SHB acknowledges support from the Georgia Tech Quantum Alliance and a Georgia Tech Seed Grant. EFD acknowledges DOE ASCR funding under the Quantum Computing Application Teams program, FWP number ERKJ347. This work used Bridges2 at the Pittsburgh Supercomputing Center (PSC) and Delta and the National Center for Supercomputing Applications (NCSA) through allocation PHY210084 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.
Data availability
The data and code supporting this study’s findings are available at https://github.com/comp-physics/fully-QLBM under the MIT license.
Conflict of interest
The authors have no conflicts to disclose.
@@ Appendix A D1Q3 Implementation
The implementations of the QLBM mentioned in section II and section III used the D1Q2 scheme, which uses two moving particles per lattice site. This scheme shows checkerboard pathologies when using a sharp function as the initial condition because it simulates two independent sub-lattice [27]. Yepez [37] presents a cure for this, allowing only one qubit to move while the other remains stationary. However, this requires two collision and streaming procedures per time step. Our results show that the quantum D1Q3 scheme, which adds a stationary qubit to each lattice site, solves this problem. A D1Q3 scheme has three particle distributions across 3 links: left, right, and stationary. So, some fictitious particles remain stationary in a time-step, smoothing the curve. Figure 9 compares the quantum solution of the diffusion equation using the D1Q2 scheme with the D1Q3 scheme for a delta function as the initial condition. These simulations were performed using the Qiskit matrix product state simulator, $45$ qubits for the D1Q3 scheme, and the sampling procedure used $10^{6}$ shots. The spatial domain consists of $N_{s}=15$ lattice sites, and results are shown after $10$ time steps. The result shows that the quantum D1Q3 solution agrees with the classical one and that the D1Q3 version remedies the checkerboard effect. We note that the D1Q3 scheme does not affect algorithmic complexity and can employ measurement avoidance analogous to the D1Q2 scheme presented in the main manuscript body.
Figure 9: QLBM solutions to the diffusion equation using the D1Q3 scheme. “Quantum - classical streaming” corresponds to the classical streaming algorithm discussed in previous work [27], “Fully quantum” refers to the present algorithm, and “Classical” is the classical implementation of the lattice methods as labeled.
@@ Appendix B Simulator Results
In fig. 10, we show the results for different quantum simulator implementations of our algorithm and its circuit. Simulators include statevector-based and tensor-network types that use the matrix product state (MPS) method [43]. Two different simulation frameworks are tested: Qiskit and XACC [44]. Statevector and MPS simulators are part of the Qiskit Aer backend. The XACC qpp (Quantum++) simulator is based on a C++ general-purpose quantum computing library [45]. XACC exatn-mps simulator is a part of the TNQVM (tensor-network quantum virtual machine) simulation backend and uses a noiseless, matrix product state (MPS) wave function decomposition for the quantum circuit [46]. The wall-clock times of fig. 7 (c) were obtained via simulation on an AMD EPYC 7742 (64-core) processor.